Vern invested $11,000. Part was invested at 15% and part at 13%. If the total interest at the end of the year is $1,610, how much did he invest at 15%?
2 pts

Answer by fractalier(6550)

Let x be the amount invested at 15%.
Then 11000 - x is the amount invested at 13%.
So we have the equation
(.15)x + (.13)(11000 - x) = 1610
Multiply all by 100 to clear decimals (makes it easier)
15x + 13(11000 - x) = 161000
15x + 143000 - 13x = 161000
2x = 18000
x = $9000, the amount at 15%
11000 - x = $2000, the amount at 13%